X Company, a manufacturer, uses a plantwide overhead allocation system with direct labor as the cost driver. The following is budgeted total cost and activity information in 2018:

Direct materials $291,758
Direct labor $510,577
Overhead $531,000
Research & development $110,000
Advertising $59,000
Direct labor hours 24,698

The following information is taken from the job cost sheet for Job A in 2018:

  Direct     Direct     Direct
Materials     Labor Labor Hours Overhead
Total   $44,770     $88,000       747       ??


How much overhead was allocated to Job A [round overhead rate(s) to two decimal places]?
Answer #1

Answer = $91,520

Plant wide overhead Rate = Overhead Labor cost = $,531,000/510,577 =104%

Overheads allocated to Job A = Labor cost * Overhead Rate = 88,000*104% =$91,520 (Answer)